Find the equation of tangent line to the circle x^2+y^2+z^2+5x-7y+2z-8=0,3x-2y+4z+3=0 and the point (-3,5,4)


1
Expert's answer
2021-02-24T06:02:19-0500

"x^2+y^2+z^2+5x-7y+2z-8="

"(x+\frac{5}{2})^2+(y-\frac{7}{2})^2+(z+1)^2=\frac{55}{2}"

This is sphere with center "C (-\frac{5}{2},\frac{7}{2},-1)" and radius "\sqrt{\frac{55}{2}}."

Find the equation of the tangent plane to the sphere at the point "P(-3,5,4)."

The radius vector or normal is "CP=\begin{pmatrix}
 -0.5 \\
 1.5\\
5
\end{pmatrix}."

A vector in the plane we seek is "v=\begin{pmatrix}
 x+3 \\
 y-5\\
z-4
\end{pmatrix}."

So, "\begin{pmatrix}
 -0.5 \\
 1.5\\
5
\end{pmatrix}*\begin{pmatrix}
 x+3 \\
 y-5\\
z-4
\end{pmatrix}=0\Rarr"

"-0.5(x+3)+1.5(y-5)+5(z-4)=0".

Answer:

The equation of the tangent plane is

"1.5y-0.5x+5z-29=0."

____________________________________________________________________________________________